1. In 20x8, BT Inc. granted a nonqualified stock option to Ms..P to buy 500 shares of BT stock at $20 per share for 10 years. At date of grant, BT stock was trading on NASDAQ for $20 per share. In the current year, Ms. P exercised the option when BT’s stock was trading at $37.10.

a) How much income must Ms. P recognize in 20x8 and in the current year because of the option?

b) What are the tax consequences of the stock option to BT Inc. in 20x8? In the current year?
Answer #1

37.10-20= 17.10 *17.10*500= 8,550 (P will not recognize any income. No Income in the year of grant and 8,550 of ordinary Inco